The pharmacy received 1000 bottles of mineral water. The probability that the bottle




will be broken during transportation is 0.003. Determine the likelihood that the pharmacy will




receive broken bottles: a) Two b) Less than two c) More than two d) At least one

1
Expert's answer
2022-01-19T16:32:35-0500

The normal approximation to the binomial distribution:

"\mu=np=1000*0.003=3."

"\sigma=\sqrt{np(1-p)}=\sqrt{1000*0.003*(1-0.003)}=1.73."


a) "P(1.5<X<2.5)=P(\frac{1.5-3}{1.73}<Z<\frac{2.5-3}{1.73})=P(-0.87<Z<-0.29)="

"=P(Z<-0.29)-P(Z<-0.87)=P(Z<0.87)-P(Z<0.29)=0.1937."


b) "P(X<1.5)=P(Z<\frac{1.5-3}{1.73})=P(Z<-0.29)=0.3859."


c) "P(X>2.5)=P(Z>\frac{2.5-3}{1.73})=P(Z>-0.87)=0.8078."


d) "P(X>0.5)=P(Z>\frac{0.5-3}{1.73})=P(Z>-1.45)=0.9265."